a jar has a mixture of hydrogen and oxygen in the ratio 1:5 
the ratio of their mean kinetic energy is
 

Grade:12

3 Answers

Arun
6 years ago
Dear Vaibhav
Mean K.E. = (3/2 )* K*T
Hence it depends only on temperature.
Thus ratio of their mean K.E. = 1:1
Regards
Arun (askIITians forum expert)
SR Roy
5 years ago
 
Mean K.E. = (3/2 )* K*THence it depends only on temperature.Thus ratio of their mean K.E. = 1:1
This is the answer
 
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
The Mean K.E. of a gas = 3/2 x K x T   ( where, K = R/NA = const.)
Hence it depends only on temperature.
Thus ratio of their mean K.E. = 1:1
 
Thanks and regards,
Kushagra